solve by elimination method
4x+3/5-2-3y/4=x+y
5x/6+ 3y/4=2x+4y+3

Answer by Alan3354(69443)

solve by elimination method
4x+3/5-2-3y/4=x+y
5x/6+ 3y/4=2x+4y+3
--------------
4x - 7/5 - 3y/4 = x + y
-7x/6 - 13y/4 = 3
-------------
3x - 7y/4 = 7/5
-14x - 39y = 36
---------------
12x - 21y = 28/5
14x + 39y = -36
-------------
84x - 147y = 196/5
84x + 234y = -216
---------------------- Subtract
-381y = 1276/5
y = -1276/1905
======================
12x - 21y = 28/5
12x - 21*(-1276/1905) = 28/5
12x + 26796/1905 = 10668/1905
12x = -16128/1905 = -1792/127
x = -1792/1524
x = -448/381